Polaski Company manufactures and sells a single product called a Ret. Operating at capacity, the company can produce and sell 34,000 Rets per year. Costs associated with this level of production and sales are given below:

  

Unit

Total

Direct materials

$

20

$

680,000

Direct labor

6

204,000

Variable manufacturing overhead

3

102,000

Fixed manufacturing overhead

5

170,000

Variable selling expense

4

136,000

Fixed selling expense

6

204,000

Total cost

$

44

$

1,496,000

The Rets normally sell for $49 each. Fixed manufacturing overhead is $170,000 per year within the range of 26,000 through 34,000 Rets per year.

Required:

1. Assume that due to a recession, Polaski Company expects to sell only 26,000 Rets through regular channels next year. A large retail chain has offered to purchase 8,000 Rets if Polaski is willing to accept a 16% discount off the regular price. There would be no sales commissions on this order; thus, variable selling expenses would be slashed by 75%. However, Polaski Company would have to purchase a special machine to engrave the retail chain’s name on the 8,000 units. This machine would cost $16,000. Polaski Company has no assurance that the retail chain will purchase additional units in the future. What is the financial advantage (disadvantage) of accepting the special order? (Round your intermediate calculations to 2 decimal places.)

2. done

3. Assume the same situation as described in (2) above, except that the company expects to sell 34,000 Rets through regular channels next year. Thus, accepting the U.S. Army’s order would require giving up regular sales of 8,000 Rets. Given this new information, what is the financial advantage (disadvantage) of accepting the U.S. Army's special order?
Answer #1

1.

Sales Revenue $        3,29,280 =8000*49*(100%-16%)
Less Costs
Direct Material $        1,60,000 =8000*20
Direct Labor $           48,000 =8000*6
Variable overhead $           24,000 =8000*3
Variable Selling Expense $             8,000 =8000*4*25%
Engraving machine $           16,000
Total Expenses $        2,56,000
Financial Advantage (Disadvantage) $           73,280

3.

Sales Revenue $        3,29,280 =8000*49*(100%-16%)
Less Costs
Direct Material $        1,60,000 =8000*20
Direct Labor $           48,000 =8000*6
Variable overhead $           24,000 =8000*3
Variable Selling Expense $             8,000 =8000*4*25%
Engraving machine $           16,000
Total Expenses $        2,56,000
Financial Advantage (Disadvantage) $           73,280
Regular Sales foregone $         -62,720 =-8000*49*16%
Variable Selling Expense Saved $           24,000 =8000*4*75%
Net Financial Advantage $           34,560